OpenStudy (maheshmeghwal9):

Calculate the wavelength transition in the Balmer series of atomic hydrogen. Answer: - 1.52300 * (10^6) (m^(-1)) = 1 523 000 m^-1

OpenStudy (maheshmeghwal9):

@Zarkon @Yahoo! @FoolAroundMath @mathslover Please help:)

OpenStudy (anonymous):

1.52300 * (10^6) (m^(-1)) = 1 523 000 m^-1

OpenStudy (maheshmeghwal9):

I want how to calculate it?

OpenStudy (anonymous):

@maheshmeghwal9 see my calculation plzz

OpenStudy (maheshmeghwal9):

but where is that?

OpenStudy (anonymous):

1.52300 * (10^6) (m^(-1)) = 1 523 000 m^-1

OpenStudy (maheshmeghwal9):

this is the answer of this question I don't want this I want solution of this question ending with this answer.

OpenStudy (maheshmeghwal9):

@Rohangrr Please see the question now:)

OpenStudy (maheshmeghwal9):

this is the actual question.

OpenStudy (anonymous):

editing the question hmm...

OpenStudy (maheshmeghwal9):

ya

OpenStudy (anonymous):

better go to this @maheshmeghwal9

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Balmer_series

OpenStudy (maheshmeghwal9):

I have all the theories but how to do this question? the link doesn't help becoz it says same as my book. no difference:/

OpenStudy (anonymous):

The wavelengths for the Balmer series are all between 410 and 656 nm. So, I'm not sure what your answer is for, frequency? Okay, If you take 1/(your answer) you get the wavelength for the first Balmer transition from the Rydberg formula.

OpenStudy (maheshmeghwal9):

the same prob is with me that no \[n_1 \space ; \space n_2\]& Z = atomic number are given therefore i can't apply this formula \[\frac{1}{\lambda}=R \times Z^2 \times [\frac{1}{n_1^2}-\frac{1}{n_2^2}].\]

OpenStudy (anonymous):

n1 = 2, n2 = 3

OpenStudy (maheshmeghwal9):

how did u gt that? wt is Z?

OpenStudy (maheshmeghwal9):

oops sorry Z=1 given in the question

OpenStudy (maheshmeghwal9):

but how did u gt this? n1 = 2, n2 = 3

OpenStudy (anonymous):

That's part of the definition of the Balmer series. All transitions end with the second orbital level: n = 2. The transition that has the wavelength for your answer starts from the third orbital level, so n2 = 3.

OpenStudy (maheshmeghwal9):

n2 may be greater than 3 too.

OpenStudy (maheshmeghwal9):

??????????

OpenStudy (anonymous):

Yes. Look at the picture of the levels, under the Physics topic, here: http://en.wikipedia.org/wiki/Hydrogen_spectral_series

OpenStudy (maheshmeghwal9):

Then why did u choose only n2=3 why not n2>3

OpenStudy (anonymous):

Because if you flip your answer: 1/(your answer) you get the wavelength from the transfer n2 = 3 to n1 = 2.

OpenStudy (maheshmeghwal9):

but if i didn't know my answer then wht did u do?

OpenStudy (anonymous):

Is the question giving you energy levels and asking for the wavelength? I've had questions that ask for the wavelengths for all possible transitions. What is the question asking for?

OpenStudy (maheshmeghwal9):

The question is only this much. Calculate the wavelength transition in the Balmer series of atomic hydrogen. & one more thing i think if u flip my answer u will get wave number not wavelength becoz my answer is wavelength.

OpenStudy (maheshmeghwal9):

becoz \[\bar \nu = \frac{1}{\lambda}\]

OpenStudy (anonymous):

I'm not sure what you mean by wave number. Your answer is the frequency for the first energy transition. So, if you flip it you get the wavelength. There are wavelengths that result from a transition from n = 4 to n = 3, and from n = 5 to n = 2 ... there are many transitions and wavelengths possible. If you just want one wavelength, I'd give the one we've been talking about, from n = 3 to n = 2.

OpenStudy (maheshmeghwal9):

ok i will see it thanx for help:)

OpenStudy (anonymous):

welcome